.237g of NaOH(s) is added to 2 moles of liquid water in a constant pressure calorimeter. The water in the calorimeter increases from 25 degrees Celcius to 43.2 degrees Celsius. What is the q of the reaction? Water has a specific heat capacity of 4.184 J/g degrees Celsius. Is this reaction exothermic or endothermic?
Answer #1

molar mass of water = 18.01528 g/mole

gm of compound = no. of mole X molar mass

gm of water = 2 X 18.01528 = 36.03056 gm

total mass of solution = 0.237 + 36.03056 = 36.26756 gm

specific heat capacity = 4.184 J/g 0C

\DeltaT = 43.2 - 25 = 18.20C

q = mass of water X specific heat of water X \DeltaT

= 36.26756 X 4.184 X 18.2

= 2761.73 J

q = 2661.73 J

tempreture is incresed in this reaction therefore heat is evolved hense reaction is exothermic